OpenStudy (anonymous):

A disc jockey must play 14 commercials during 1 hr of a radio show, each commercial is either 30 seconds or 60 seconds long if the total commercial time durring 1 hr is 11 min how many 30 sec commercials were played that hour?

OpenStudy (nikita2):

x*30 + y*60 = 11*60 sec. x+y=14

OpenStudy (nikita2):

x = 6, y = 8

OpenStudy (anonymous):

so how many 30 sec commercials

OpenStudy (nikita2):

6

OpenStudy (anonymous):

what about 60 sec com

OpenStudy (nikita2):

8

OpenStudy (anonymous):

thank you can you help me with just a couple more please ?

OpenStudy (nikita2):

if i could )

OpenStudy (anonymous):

ok thank u :)

OpenStudy (anonymous):

Snookers lumber can convert logds into either lumber or plywood in a given day the mill turns out twice as many units of plywood as lumber it makes a profit of $25 on a unit of lumber and $10 on a unit of plywood how many of each unit must be produced and sold in order to make a profit of 12750?

OpenStudy (nikita2):

x*10 + y*25 = 12750 x=2y http://www.wolframalpha.com/input/?i=x*10+%2B+y*25+%3D+12750%2C+x%3D2y so x = 1700/3, y = 850/3

OpenStudy (anonymous):

SO HOW MUCH PLYWOOD AND HOW MUCH LUMBER

OpenStudy (nikita2):

Oh, x is amount of PLYWOOD and y is amount of LUMBER

OpenStudy (anonymous):

ok ty so much can u help me with a few more

OpenStudy (nikita2):

let's try)

OpenStudy (anonymous):

snookers lumber can convert logs into either lumber or plywood in a given day the mill turns out twice as many units of plywood as lumber it makes a profit of 30 dollars on a unit of lumber and 45 $ on a unit of plywood how many of each unit of lumber and units of plywood to make a profit of 12240?

OpenStudy (anonymous):

another one im having a hard time with are 5x+5y=13 7x+89=y use substitution method

OpenStudy (anonymous):

the other one is 8x-9y=24.5 7y-2x=8.5 what is the solution

OpenStudy (anonymous):

I think nikita deserves a couple of medal here. :)

OpenStudy (anonymous):

5x+5y=13 7x-2y+19 what is the solution type an ordered pair or integer or fraction simplify for all

OpenStudy (anonymous):

5x-6y=16 7x+89=y what is the solution to the ordered pair use substitution method

OpenStudy (nikita2):

x + y = 13/5 7x - y = 89 x + y = 13/5 6x + 0 = 89 -13/5 => x = (89 - 13/5)/6 y = 13/5 - x => y = 13/5 - (89 - 13/5)/6

OpenStudy (anonymous):

is this for the word problems

OpenStudy (nikita2):

it is for another one im having a hard time with are 5x+5y=13 7x+89=y use substitution method

OpenStudy (nikita2):

for word problem : x*45 + y*30 = 12240 x=2y and x = 204, y = 102

OpenStudy (anonymous):

ok so 204 stands for what and 102 is what

OpenStudy (nikita2):

x is amount of PLYWOOD and y is amount of LUMBER

OpenStudy (anonymous):

ty :)